### v1

# Identifying the post-pandemic determinants of low performing students in Latin America through interpretable Machine Learning SHAP Values—Insights from PISA 2022

Abstract
The high prevalence of students not achieving the basic competencies in Latin America is concerning. Even more so given the region’s deep structural inequalities and the larger post-pandemic regional learning losses. Within this scenario, this paper contributes to the identification of the determinants of bottom and low performers (below level 2) using recent advancements on explainable machine learning methods. In particular, relying on PISA 2022 data for 10 countries and using the Shapley Additive Explanations (SHAP) analysis, I identify critical factors impacting on the student performance across low performers groups. I find that a student with the highest probability of being a not achiever speaks a minority language and had repeated, has no digital devices at home, comes from a poor family and works for payment half of the week, and the school he//she attends has wide disadvantages such as bad school climate, weak ICT infrastructure and poor teaching quality (only a third of teachers being certified). Regarding countries’ estimates, I find quite homogeneous patterns as far as global average contribution of top ranked factors is concerned, with repetition at primary, household wealth, and educational ICT inputs being top ten ranked covariates in at least 8 out of the 10 total countries. The paper findings contribute to the broad literature on strategies to identify and to target those most left behind in Latin American education systems.
Keywords learning poverty $\cdot$ interpretable machine learning $\cdot$ SHAP values $\cdot$ SDG4 $\cdot$ Latin America $\cdot$ PISA 2022 $\cdot$ COVID-19 pandemic
1 Introduction
Latin American students’ poor achievement at secondary level (Arias Ortiz et al., 2024b) and weak learning growth (Angrist et al., 2021) are of great policy concern, even more so because the region’s low learning trap is rooted in systemic and profound inequalities (Breton and Canavire-Bacarreza, 2018; Fernández et al., 2024). Based on the OECD Programme for International Student Assessment (PISA) 2022 data, Arias Ortiz et al. (2023) estimate that 75% of the students in the region are low math performers —not reaching the basic competencies given by level 2—, and more than half of students (around 55%) lacks basic reading skills (fail to understand a simple text). Other recent regional learning survey results such the Regional Comparative and Explanatory Study (ERCE 2019) also points in the same direction on the extent of the region’s learning crisis for primary school students (UNICEF et al., 2021).
This educational crisis has been aggravated post-pandemic (Azevedo et al., 2021; Betthäuser et al., 2023; Bracco et al., 2025; Neidhöfer et al., 2021) because the Latin America (henceforth: LAC) region experienced the longest period of school closures with poor connectivity (Gómez and Andrés Uzín P, 2022; Munguia, 2023), where, on average (between 2020 and 2021) schools were closed for 269 days; representing a disruption of 1.42 years (Bracco et al., 2025). Hence, the pandemic pushed a large cohort of students down in the learning distribution affecting the more disadvantaged, with estimates suggesting the LAC region lost around 0.9-1.1 years of schooling (Azevedo et al., 2021). In the case of OECD countries the share of disadvantaged low performers rose by 9% in the 2018-2022 period (OECD, 2023c), whereas in Latin America (already suffering from long term broad educational inequalities Brunori et al., 2024; De la Torre et al., 2017; Delprato et al., 2015) according to Delprato and Antequera (2025), using PISA 2022, the proportion of low performers (below level 2) rose by 5% in math and slightly declined for reading, and remained stable in science (Arias Ortiz et al., 2023). Yet, good cognitive skills are fundamental for quality jobs and to enhance future labour prospects of secondary students which are vital for a steady growth of the economies in the region (Angrist et al., 2021; Hanushek and Woessmann, 2023). It is within this post-pandemic scenario, that the paper addresses the drivers of learning poverty in the LAC region.
The COVID-19 pandemic worsened the already prevailing and detrimental educational inequality in the LAC region. Prolonged school closures led to loss of learning, increased inequalities, having negative impacts on psychosocial health and well-being of students and teachers, with younger children and students from more vulnerable groups experiencing the greatest losses (Asadullah et al., 2023; UNESCO, 2024). For instance, Bracco et al. (2025)’s simulations show how educational loss is tightly related to households’ socioeconomic levels. The authors estimate a mean loss years of education for a student in the bottom decile of the income distribution of 81%, nearly four as large than a student in the top decile (losing 22%).
Household’s income shocks leading to increasing poverty (which in the region are estimated between 23%-39%; Bracco et al., 2025) due to the pandemic and their consequences on educational outputs compounded established drivers of learning poverty in LAC region. Established drivers of low performance in the region are repetition (Randall and Anderson, 2016; Vinas-Forcade et al., 2021), parental education and Indigeneity (Cortina, 2017; Delprato, 2019), educational inputs (e.g., households’ ICT and books stocks, family educational support: Lara and Saracostti, 2019; Torrecilla and Hernández-Castilla, 2020), as well as school characteristics such as school contextual poverty, school type and poor school climate (Berkowitz et al., 2017), ICT investments (number of PCs per student, rate of connectivity; Agasisti et al., 2023), quality of teaching —proxied by teachers uptake of professional development//PD and qualifications (McGinn and Schiefelbein, 2023).
Yet, these drivers are not likely to be homogenous across sub-groups at the bottom of students’ learning distribution, and there is a need of carefully distinguishing these groups for targeting. So, in this paper, I fill this gap by investigating how these determinants may vary along low performing groups in the region. Importantly, this empirical contribution is carried out using the latest PISA available data for the LAC region after the COVID-19 pandemic, offering the latest results on students’ low learning levels determinants. Furthermore, studies using interpretable machine learning (ML) methods based on learning surveys addressing determinants of students’ achievement are rather scarce (e.g., Bu and Chen, 2023; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022), with their focus being on richer countries education systems whose realities are entirely contrasting to global south education systems. Methodologically, I contribute here to this small body of literature, expanding it to global south education systems, and at a regional scale.
In particular, relying on the latest post pandemic (year 2022) PISA data for 10 LAC countries, and recent developments in the field of explainable machine learning: Shapley Additive Explanations (SHAP) analysis (Chen et al., 2023; Lundberg and Lee, 2017), this paper attempts to answer the following research questions for students at the bottom of the learning distribution (from level 0 to level 2 achievement groups):
  1. 1.
    Globally, which are the principal determinants of low performance for the whole LAC region?
  2. 2.
    What are the learning profiles (namely: leading determinants) of students with the highest and lowest chances of poor achievement across the region? This question’s approach has a local dimension as it deals with specific students (at extremes).
  3. 3.
    Which are the main drivers at the country level, both globally and locally, of learning poverty?
The rest of the paper is organised as follows. Section 2 contains a review of ML explanatory studies within the educational research field. In Section 3 I describe the data used, the definition of the working samples for the two educational outcomes and leading covariates used. Section 4 describes the methodological approach —i.e., the array of learners (models) and Shap values methodology for interpretability of results and the analytical steps followed. Findings are presented in Section 5, both at the regional level and country level. In Section 6 I discuss the implication of results for future targeting of low performing students in the region and, in Section 7, I include some concluding remarks.
2 Background—machine learning and educational studies
Underpinned by their complexity, machine learning (ML) models are now widely applied to solve supervised learning problems in many different fields and settings (Olsen et al., 2024; Chen et al., 2023; Dhal and Azad, 2022; Sarker, 2021) because supervised ML methods are capable of generating more precise predictions than standard statistical models (Chen et al., 2022; Lisboa et al., 2023). Yet, this may come at the expense of model interpretability (Ruding, 2019). Therefore, to account for this absence of interpretation (Olsen et al., 2024), the field of explainable interpretable machine learning has become an area of recent substantial research (Covert et al., 2021; Linardatos et al., 2020; Holzinger et al., 2020; Molnar, 2020).
In tandem with this methodological progress, there has also been a recent growth the application of ML, broadly in the social sciences (Brand et al., 2023; Grimmer et al., 2021) and within the educational field, which is still limited, but continuously grows (e.g., Ersozlu et al., 2024; Hilbert et al., 2021). Grimmer et al. (2021) advocate for using an agnostic approach, using machine learning to make social science inferences where new research questions are generated, and concepts and hypothesis can be tested with new data, rather than assuming theory are robust before investigating what is emerging from available data. Brand et al. (2023) describe how to use ML methods related to causality, mostly interest in learning about heterogeneous treatment effects and to minimize biases in estimating population-level quantities of interest (also see the review of Athey and Imbens, 2019 for the economics and econometrics fields).
Hilbert et al. (2021) show how the robustness of educational research can benefit from ML methods and the actual prediction of new data and validation of empirical models, specially when the size of data is large like in international learning surveys (e.g., TIMSS, PIRLS, PISA). In the educational research field, so far, most of the focus has been on prediction, such as academic performance in regression tasks (Hellas et al., 2018; Öz et al., 2024) and, to a lesser extent, for the identification of students at risk (e.g., Alam and Mohanty, 2022; Rodríguez et al., 2023). Dol and Jawandhiya (2023), in a review of nearly 150 articles over 2010-2020 period, find out that classification techniques are mostly used technique for analysing students performance within the Educational data mining (EDM) field. Similarly, Ersozlu et al. (2024), in a systematic review of ML methods used for educational data, find that nearly 90% of publications use ML approaches for predicting students’ performances and finding learning patterns. In spite of the prediction-centred approach of ML, researchers have recently produced various approaches to better understanding of models by quantifying the impact and interactions of features (Hilbert et al., 2021). The SHAP method (Lundberg and Lee, 2017), which I employ in the current paper, is a leading example of this, falling into the group of interpretable machine learning (IML) techniques.
Yet, the majority of educational empirical work (so far, a small body of studies) using the SHAP method are either for specific richer countries or regions, without a scaffolding inductive logic of analysis following layers of education systems (Codiroli Mcmaster and Cook, 2019) and how educational inequalities are produced, specially when applied to monitor SDG4 inequalities (Delprato, 2024).
Studies using IML methods based on learning surveys addressing determinants of students’ achievement are rather scanty (Bu and Chen, 2023; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022; Zhu et al., 2025), and subject specific. These studies’ focus is on more developed countries with more resources-intensive education systems (e.g., Spain: Gómez-Talal et al., 2024, or the US: Zhu et al., 2025) or groups of countries//regions (East Asia: Bu and Chen, 2023; Huang et al., 2024) whose realities are entirely contrasting to global south education systems. The only exception is Ali et al. (2025), which analyse factors influencing academic performance among primary school students in Somaliland; though, this study, is a combination of ML and regression methods, and with the latter replacing IML techniques as Shapley values. Methodologically, I contribute here to this emerging and scanty group of studies, but expanding it to global south education systems of Latin America as well as at a regional scale.
3 Data
The analysis is based on the latest wave of the OECD PISA (Programme for International Student Assessment) for the year 2022 which measures secondary students’ learning achievement (at 15 years-old), i.e., their ability to use their reading, mathematics and science knowledge and skills to meet real-life challenges.1 The paper is focused on the LAC region, which includes a sample of 10 countries, namely: Argentina, Brazil, Chile, Colombia, Dominican Rep., Mexico, Panama, Peru, Paraguay and Uruguay.2 Since the aim of the paper is about identifying the determinants of learning poverty (i.e., those students located at the bottom of the learning distribution in three subjects: math, reading and science) in the region, I employ two comparisons: no competency achieved (level 0 which includes level1b and level1c) versus level 1 (level1a), and level 1 versus level 2 (where those below level 2 are referred to as ‘low performers’; OECD, 2023b, p. 89). These two working samples contain, respectively, between 16,236 and 9,484 observations.
It should be mentioned that the PISA 2022 data gathering was mediated by the pandemic. Originally, it was planned to take place in 2021 but the disruption caused by COVID-19 global pandemic forced the assessment to be postponed by a year, to 2022, whose content was updated with further questions linking the “resilience of education systems” from the point of view of fairness and inclusion within the pandemic’s remote learning (OECD, 2023a).
3.1 Data coding—rationale
My focus on the three bottom levels of achievement is aligned to the observed adverse long-term learning trajectory of the region where, for example, the rate of low-performers in math among the poorest students either increased or remained unchanged across all countries since 2012 (Arias Ortiz et al., 2024a, b). As said above, I use two binary educational outcomes (or dependent variables) among the low performance group which are cardinally defined towards those students achieving the least: level 0 (=1) versus level 1 (=0), and level 1 (=1) versus level 2 (=0). As shown in Table 1, these two comparisons//samples capture most of the LAC students. For instance, levels 0 to 2 (first three rows of Table 1) account for 89.5% (math), 77.5% (reading) and 80.9% (science) of the total students.
[Table 1 here]
Regarding data processing, rather than analysing low achievement drivers for specific subjects as in other PISA empirical ML studies (see: Gómez-Talal et al., 2024; Huang et al., 2024, and the systematic review: Wang et al., 2023), I follow a subject-aggregated definition of dependent variables –for instance, $y_{\text{level 0,level1}}$ = 1 if a student falls into level 0 for math, reading and science and equals to 0 if a student falls into level 1 for the three subjects. This leads to rates for the first educational outcome (levels 0 versus 1) of 65.3% and for the second educational outcome (level 1 versus level 2) of 59.3% (Panel B of Table 1). I constructed two datasets for each educational outcome to be analysed separately: a dataset with student//family variables and a dataset with school variables. Besides, after selecting leading covariates with lower missing rates, I eliminated records with missing values (e.g., $\approx$ 3% in the student//family and school datasets).3 Full details on covariates construction and definitions are shown in Appendix C (Tables C1 and C2).
3.2 Summary statistics
3.2.1 Students and family covariates
The empirical relevance of students and family factors as drivers of the probability of falling into the sub-groups of bottom performers (level 0 versus level 1) and low performers (level 1 versus level 2) can be seen by the summary statistics of Table 2. On the one hand, in comparison to students whose achievement fall into level 1, level 0 (no achieving) students (Table 2, columns 1 to 3), are likely to do less homework, have a lower number of digital devices (= -1.72) and fewer books (= -0.35), being twice as likely to be from an immigrant family and three time as likely to speak another language at home (Panel A, columns 1 to 3). Repetition for the level 0 group of students —either at primary or lower secondary— is much larger (33% against 9% and 22% against 9%) as well as the chances of missing or skipping school as the intensity of paid work (1.66 versus 1.45). Further, students from the level 0 group compared to students from level 1 group are poorer (69% comes from families in the bottom half of the family SES index distribution in comparison to 59% in the case of level 1 students) and come from larger families (Panel B, column 3, Table 2).
[Table 2 here]
On the other hand, these gaps on covariates are similar and hold when comparing level 1 versus level 2 students (Table 2, columns 4 to 6), although with some differences in magnitude with the level 0-level 1 comparison. For instance, the gap on possession of books at home and chances of paid work are larger, but repetition gaps are smaller. Notably, the gaps on parental education become relevant (and statistically significant) as well as the importance of family educational support.
3.2.2 School covariates
As shown by Table 3, school context —be its composition, inputs and infrastructure, teachers profile or policies implemented— matters when related to the chances of students falling into either the lower achieving group level 0 (among bottom performers, columns 1 to 3) or level 1 (among low performers, columns 4 to 6). The influence of school context is slightly more prominent when comparing the bottom two levels of performance. Here, I mention some key differences. When contrasting to students who achieve in level 1, level 0’s students are more likely to come from rural schools (by 13%) and public schools (by 12%), and from schools where there is less competition, which are located in smaller towns, which received less funding from fees and which are smaller in terms of enrolment (Table 3, column 3). Likewise, students at the bottom of the learning distribution attend school where there are less PC for teaching connected to the web and with less involvement of parents and with fewer methods for students’ testing.
[Table 3 here]
In addition, the school prevalence of disadvantage students now becomes an important factor between level 1 and level 2 groups, being 9% higher for the latter group (column 6, Table 3). The remaining covariate raw differences are, to a large extent, similar to the previous students’ comparison. There are some distinct pattern in covariates differences, however. For example, no statistical differences are found for STRs, the proportion of certified teachers, infrastructure, and in some school policies (e.g., transfer, improvement policies, and ability grouping). On the whole, summary statistics from Tables 2 and 3 underline who contextual disadvantages stemming from student, family and school factors are vital for the likelihood of students falling into the lowest performing groups in the LAC region.
4 Methods
4.1 Binary classification models
Given the dichotomous nature of dependent variables, I employ five binary classification ML models (for details of the five ML models below, see: Hastie et al., 2009, 2015 and for its implementation: Pedregosa et al., 2011; Raschka and Mirjalili, 2019), for the two dichotomous dependent variables (level 0-level 1 and level 1-level 2). Each model was trained using the stratified $K$-fold (5 folds) method and train-test split of 80%-20% of the data so that to ensure the data is balanced within each fold within this undersampling strategy. Each model was optimised using different hyperparameters setup and the application of grid search to find the top performing model version among each type of model. Below, I briefly describe each of the five ML models employed.
4.1.1 Logistic regression
Logistic regression (or just logit) is one the principal models for binary outcomes, which predicts the probability that the outcome equals to one conditional on a vector of covariates and the error term following a logistic distribution:
$Pr\big(y_{i}=1|\mathbf{x}\big)=\text{exp}\big({\beta_{0}+\mathbf{x}_{i}\boldsymbol{\beta}}^{\prime}\big)/\big(1+\text{exp}\big({\beta_{0}+\mathbf{x}_{i}\boldsymbol{\beta}}^{\prime}\big)$ (1)
Logit is simpler and easier to interpret. Here, the hyperparameters used for the two versions are lack of regularisation (no penalty), and with penalty L2 (Ridge) option.
4.1.2 Logistic Lasso with CV penalty
I also use a different version of logistic regression; namely, a Logistic Lasso (with cross-validation, CV) (logitlasso) which imposes a penalty (L1) on the size of coefficients to control overfitting and the size of it is chosen by CV. The lasso penalises the absolute size of coefficients via the penalty parameter $\lambda$, yielding sparse solutions where some coefficient estimates are set to exactly zero. I used four values for $C$ (the inverse of regularization parameter values) for cross-validation ($C$ = 10, 50, 100 and 200). The penalised objective function for lasso is:
$Q_{\mathcal{L}}=\sum_{i=1}^{N}\widetilde{w}_{i}\times\mathcal{L}+\lambda\sum_{j=1}^{P}\kappa_{j}|\beta_{j}|$ (2)
where $\mathcal{L}$ is the logit likelihood function.
4.1.3 Random forest
Random Forest (RF) improves modelling by ensembling a combination of multiple decision trees and then aggregating their predictions. RF trains several decision tree models (learners) separately and then let all the learners decide the output of the test samples –the final classification– by a voting mechanism. That is, the predicted class is: $\widehat{\mathbf{y}}=\text{mode}\big\{y_{1},y_{2},...,y_{n}\big\}$, and $\mathbf{y}_{i}$ is the prediction of the $i$th tree. Also, the training samples of each learner are randomly chosen and when building the decision tree a portion of the features (variable) is randomly selected from the features to build the decision tree. The hyperparameters assessed are the number of estimators (or trees), with 100, 500 and 1,000 trees chosen and two maximum depths of each tree given by 5 and 10 leaves.
4.1.4 Gradient boosting
Another ML classifier method I use is Gradient Boosting (GB) under six formats (or combination of hyperparameters): the number of trees ranging from 100, 500 and 1,000 and two learning rates (1% and 10%). GB is an ensemble of predicted decision trees models run in a sequential fashion correcting prediction of previous models, and the update rule is:
$f_{m}=f_{m-1}(\mathbf{x})+\sum_{j=1}^{J_{m}}\gamma_{jm}1(\mathbf{x}\in R_{jm})$ (3)
where $f_{m}$ is the model $m$ interaction and $J$ are terminal nodes, $\gamma_{jm}$ the fitted values and $R_{jm}$ the residuals.
4.1.5 Neural networks
Neural nets (NN) or Feed-forward neural networks consist of hidden layers that link the predictors (or input layers) to the outcome. Each hidden layer is composed of multiple units (nodes) that pass signals to the next layer using an activation function. Central tuning parameters are the choice of the activation function and the number and size of hidden layers. I consider six specifications for the NN: one neural net with one hidden layers of 200 nodes each and another neural net with three hidden layer set of 200, 100 and 50 nodes (200, 100, 50), using three activation rules.
4.2 Assessment of binary classification models
I assess the performance of each of different ML model versions (made by the configuration of hyperparameters) relying on an array of evaluation metrics (Naidu et al., 2023). Specifically, I use six metrics when comparing the models: Accuracy (ACC), Recall (RC), Precision (PR), F1 Score (F1), Specificity (SP) and the Area Under the Curve (AUC). These metrics provides overall measures of model performance. Noteworthy is the AUC indicator because, by calculating the area of under the Receiver Operating Characteristic (ROC) curve, it can efficiently balance the true positive rate and the false positive rate. I include a description of the six metrics in Appendix 0.A.
4.3 Stacking approach for classification
When faced with a classification task as in the current paper (and after choosing the best model based on evaluation metrics) the next decision is to assess which one is best suited. But this not obvious, and combining multiple learners (ML algorithms) into a final ensemble model, is likely to lead to superior performance compared with each individual learner. Stacked generalisation (or stacking), a concept introduced by Wolpert (1992), allows a form of model averaging using a technique that utilises the predictions of various models to train a meta-model. Theoretically, stacking perform as well as individual learners (Van der Laan et al., 2007). Here, I follow the stacking approach among the best five binary classification ML models using the work of Ahrens et al. (2023), implementing stacking in Stata.
The stacking process for binary classification exercises consists of two steps. First, cross-validated predicted probabilities for each of the five base learners and each observation $\big(\widehat{y}^{-k(i)}_{(j),i}\big)$ are obtained for all folds ($K$ data partitions of similar size) expect from the fold where current observation is located (i.e., k-k(ii)). Second, the final learner (stacked model) is fit by using the observed $y_{i}$ outcome and cross-validated predicted probabilities as predictors $\big(\widehat{y}^{-k(i)}_{(1),i},\widehat{y}^{-k(i)}_{(2),i},...,\widehat{y}^{-k(i)}_{(J),i}\big)$, with the condition that weights are zero or positive ($w_{i}\geq$ 0) and the sum of weights are equal to one $\sum_{j=1}^{J=5}w_{j}=1$, where $j$ denotes different base learners. That is, for the standard least squares:
$min_{w_{1},...,w_{J}}\sum_{i=1}^{N}\Big(y_{i}-\underbrace{\sum_{j=1}^{J}w_{j}\widehat{y}^{-k(i)}_{(j),i}}_{\widehat{y}_{i}^{*}}\Big)^{2}$ (4)
where $\widehat{y}_{i}^{*}$ are the stacking predicted probabilities, a weighted sum of predicted probabilities from rerunning learner $j$ on the full sample $I$.
4.4 SHAP values—interpretable ML estimates
In the body of educational research using ML (for a review, see: Ersozlu et al., 2024; Hilbert et al., 2021), there has been a latest growth in the application of Shapley values to identify main features behind educational performance in educational surveys (for instance, Gómez-Talal et al., 2024; Huang et al., 2024). Here I define the key concepts related to Shapley values (a more detailed and technical description of Shapley values can be found in Appendix 0.B). Importantly, I rely on SHAP values to present leading estimates in Section 5.
Shapley values have their origin in game theory (Shapley, 1953), measuring the average marginal contribution of a player in a cooperative game. In the context of ML, the player is interpreted as a characteristic (or attribute), and the cooperative game becomes the prediction task performed by the model. Through recent developments (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017), Shapley values have been adapted as an explicative framework for ML models. For the estimation of the main determinants of learning poverty in the LAC region, in particular, I follow the SHAP method (Lundberg and Lee, 2017). Plainly, the SHAP method locally explains a model output by calculating the degree to which each feature (or variable) contributes to a given prediction value. The prediction for the $i$th sample is:
$\widehat{y}_{i}=\widehat{y}_{\text{base}}+\widehat{y}_{i1}+\widehat{y}_{i2}+...+\widehat{y}_{iJ}$ (5)
and where $\widehat{y}_{\text{base}}$ is the average of predictions across all $i$ samples, $\widehat{y}_{ij}$ is the SHAP value of the $j$ feature ($j$ = 1,…,J; with $J$ being total number of feature) in the $i$th sample measuring how much feature $j$ contribute to the prediction of sample $i$. Positive SHAP values ($\widehat{y}_{ij}>0$) indicates that the given feature $j$ has a positive impact on the predicted value of sample $i$, and vice versa.
4.5 Analytical steps
The steps for the analysis for the PISA 2022 data are displayed in Figure 1. First —and after carefully processing the PISA data and generating each dataset for the dependent variables— the dataset is divided into a trained dataset (80%) and a test dataset (20%). This split, coupled with the stratified $K$-fold (number of folds = 5) cross-validation, ensures that each fold retains proportionality for $y$ educational outcomes and covariates avoiding skewed results. Second, using the train dataset, I run the hyperparameters search fitting different configurations for the five ML classification models, and obtaining their best ones based on the six evaluation metrics. This yields the stacking model based on the train sample. Third, I re-estimate the chosen models excluding a model if it is weight during stacking is equal to zero (interaction 2 in Figure 1) obtaining the SHAP values. Fourth, the final stacked train data model SHAP estimation is then applied to the test data leading to main features covariates selection globally. In the final step, I produce a local version of SHAP values by examining the profile of leading explanatory variables (covariates//features) for students at the extremes of the distribution of the aggregated SHAP values, i.e., $\sum_{j=1}^{J}\phi^{(i)}_{j}$ equal to the minimum and maximum across all $i$ samples (observations).
[Figure 1 here]
5 Results
This section presents the findings via the SHAP analysis for the two samples at the bottom of the learning distribution: level 0 (no achievement) versus level 1, and level 1 versus level 2, both globally (across all samples) and locally (bottom and high performers students within each working sample), dividing the explanatory variables into two sets: students and family covariates, and school covariates. Before, I assess the five binary classification models through evaluation metrics. Countries estimates are also shown as local explanations for specific students based on the SHAP contributions index.
5.1 Models’ comparison
I present a comprehensive comparison of the efficacy of the five models using a range of evaluation metrics of Section 4.2 for the two working samples. I use a stratified KK-5 fold method approach with 80% of the PISA data working sample belonging to the trained data and the remaining 20% to the test data. The metrics for all models’ configurations are shown in Tables D1 and D2 (Appendix D). For the level0-level1 comparison dataset (the bottom two performing groups), less parsimonious models outperform other simpler models in terms of the degree of accuracy of the classification of students into achieving levels. Notably, GB and RF models consistently show being more effective in terms of the SP and AUC (columns 5 and 6, Table D1), either using more trees and depth and higher learning rates. Results for the level1-level2 comparison dataset (low performing group’s comparison; Table D2), follow an almost identical pattern, with the GB model in particular showing metrics as a leading model.
Table 4 presents evaluation metrics values for the selected better performing models (columns 1 to 6) and each model’s weight (column 7) for the stacked (ensemble) modelling approach where, in the first interaction, it includes all learners//models. Vitally, the stacking model has the benefit of combining the strengths of the individual models it comprises. Figure 2 displays the AUC for all models, including the stacking model where one can see how the classification and identification of the binary categories of educational outcomes improve by the amalgamation of models.
[Table 4 and Figure 2 here]
In the first interaction, Table 4 shows that models (learners) with zero weight —i.e., not contributing to the overall prediction of the stacking model— are logit (for the level0-level1 comparison, Panel A, column 7), and logit and neural net (for the level1-level2 comparison; Panel B, column 7). These models are therefore excluded in the second interaction. This leads to the following two stacking models used for predictions and derivation of Shapley values: lassologit, GB, RF and NN (level0-level1 data), and lassologit, GB and RF (level1-level2 data). Also, note that the weights of the final stacking model suggest the relevance in the overall prediction of GB and lassologit models.
5.2 Whole sample—global explanations. Student, family and school determinants
[…]
On the one hand, focusing on the top ten most impactful student//househol explanatory variables with the highest SHAP average values (columns 1 and 3 of Table 5), there is a similarity on determinants for bottom and low performers (with 7 out of 10 covariates being in the top ten for these two groups). Primary repetition, sense of belonging to school and wellbeing support from teachers are key drivers of falling into the lowest achievement group among bottom performers (level 0, no competency, column 1), but not in low performing group comparison (column 3). In the reverse direction, engagement in paid work, the intensity of homework and mother’s educational level are among the top ten SHAP values for level 1 versus level 2 comparison, although not for the bottom two levels of achievement comparison. Furthermore, there are some discrepancies on the SHAP values ranking among the two samples. For instance, family SES and family educational support are the leading top two determinants for low performers; yet, in the case of bottom performers sample, they are ranked fifth and sixth in order of importance. Likewise, the number of digital devices and school climate (ranked second and third) are more relevant for bottom performers, and gender having a larger impact when comparing level 1 versus level 2 achievement. Hence, this suggests distinct factors for those students at risk of falling into either level 0 or level 1 of each sample, and therefore the necessity of different strategies when identifying and targeting those students.
[Table 5 here]
On the other hand, as regards to school characteristics SHAP values (Table 5, columns 2 and 4), school size is equal important for both samples, though the rate of school disadvantage and the STR for pedagogic support are higher in the ranking for low performers in comparison to bottom performers. Notably, school type and government funding are more powerful school features for the low performing sample (column 4) than for the bottom performing sample (column 2), whereas the rate of students with a second language and wider support for inclusivity are more relevant for the latter comparison (level 0 versus level 1). In both groups of samples, policies (ability grouping, transfer, admission) are located at the lower end within the SHAP ranked values.4
Figure 3 displays the distribution of SHAP values for each student and covariate for the bottom performing group using beeswarm (top figures) and heatmap (bottom figures) plots5 which offer a more granular approach than the aggregated SHAP values of Table 5 in terms of impacts on predictions.6 All variables are shown in the order of global feature importance, the first one being the most important and the last being the least important one. Below, I explain some of the main findings emerging from these plots.
[Figure 3 here]
Figure 3’s beeswarm plot (on the left) shows how high values (displayed in red) of repetition at primary (=1) have a high positive contribution on the prediction (that is, the probability of student’s achieving level 0 instead of level 1), and an equivalent pattern is found for repetition at lower secondary, speaking other language and for those who started primary late and female students. Most of the number of devices’ SHAP values are low and fairly spread (shown in blue), and they have a positive contribution on the chances of students’ falling into level 0 achievement as well as family SES (poorer students) and any engagement in paid work (mostly blue dots), while for teacher wellbeing support estimates are equally divided between high (red) and low (blue) values, something akin is found too for number of books. This is complemented by the SHAP values heatmap for student//family covariates (bottom left) where one can see that covariates with significant homogenous segments (mostly in red) are repetition, devices number, school climate and safety and minority language spoken at home.
For school’s variables (plots on the right of Figure 3) the beeswarm plot shows that, school’s features specific estimates withholding students moving from level 1 to level 2 (with the majority of SHAP values located in the positive section of the x-axis), are low values of testing methods of students (or no testing), high students’ contextual disadvantage, low values of either the ICT index (or weaker ICT infrastructure) or STR (pedagogic staff). And, in the oppositive direction, high values (shown in red) of autonomy, private school and testing, rate of certified teachers and teachers PD attendance, for instance, can all lead to negative predictions of the model output (or higher chances of reaching level 2 achievement across the three subjects). The heatmap for school covariates, in turn, shows that covariates with darker red (with positive impacts on prediction) or blue (negative impacts on prediction) compact segments are: testing, school size (enrolment) and type, and ICT infrastructure.
5.3 Experiment—local explanations. Students at extremes of SHAP index distribution
In this section, I move away from global explanations, shifting the focus towards local explanations of SHAP values which are given by specific observations. This exercise is very relevant policy-wise. For this, I proceed as follows. First, I obtain the overall average contribution of SHAP values across the sample for each student, namely: $\Phi_{i}=\frac{1}{J}\sum_{j=1}^{J}\phi_{j}^{(i)}$; for $i$ = 1,…,$N$. Secondly, I sort out these values and, finally, I select students with the minimum ($\Phi^{\text{min}}_{i}$) and maximum ($\Phi^{\text{max}}_{i}$) contributions so that to profile the impact of covariates at extremes. Once student pairs are identified, I plot leading covariates impacts for each —and fundamentally— with plots showing the specific value each covariate takes. This exercise is important insofar as profiling learning drivers of students and their values7 and, by choosing students at extremes of the SHAP contributions, it is possible to identify students who are the most likely to be performing at level 0 or level 1 (or at level 1 and level 2 for the low performing sample) and their related covariates pushing students’ performance towards either of them.
5.3.1 Level 0 versus level 1—estimates
Here, I compare the profile of students ($\Phi^{\text{max}}_{i}$ = +7.49 and $\Phi^{\text{min}}_{i}$ = -5.92) for the bottom performing sample. In the top two plots of Figure 4 are SHAP values for the student//family group of covariates. A comparison of Figure 4a and Figure 4b demonstrates how inequality operates at the bottom of the learning distribution. A student with the highest probability of being a level 0 achiever attends a school with bad school climate (index of -0.65 falling into high interval); speak other minority language; has no digital devices; has a very weak sense of belonging to school; repeated in primary; and comes from a very poor family (SES index of -2.33 falls into the bottom half of SES distribution); and is engaged in unpaid work (1 day) and paid work (=2.5; between 2 and 3 days per week) (see: Figure 4a). In the oppositive case, a student whose performance falls into level 1, his//her family is much better off (SES index falling into top half); has books (=3) and high number of devices (=10); is satisfied with life and supported by teachers wellbeing wise; have not repeated a grade and does not work; and comes from smaller families (1 sibling) (Figure 4b).
[Figure 4 here]
School covariates, too, are dissimilar for these two students’ profiles (Figure 4c and Figure 4d). Consider the first ten ranked covariates in terms of impacts. Smaller schools (enrolment = 129), with nolangtest variable equals to 1 (hence an Indigenous school) and a huge degree of disadvantage (98%) and with weaker support for inclusivity as well as ICT, no PCs and only a third of teachers certified, are all features of schools linked to level 0 achievement. On the contrary, private schools with low STR for pedologic staff (20.2 staff per student versus 109.5 per student for the comparison), low contextual disadvantage and higher autonomy (index = +1.27) and nearly all PCs connected to the web (90.9%) and nearly 55% more qualified teachers, are all school features pushing up the chances of a bottom performer towards a level 1 achievement.
5.3.2 Level 1 versus level 2—estimates
I repeat the same exercise here for the low performer sample (level 1 and level 2). Results are displayed in Figure 5. First, compared to the student with the highest chance of reaching no more than level 1 (Figures 5a and 5b), the student reaching level 2 with the lowest index ($\Phi^{\text{min}}_{i}$) have larger family educational support (index = +1.11); 4 books (instead of 1) and 12 digital devices (instead of 6); a high index regarding school belonging (+2.37 versus -4.47); not engage in paid work and come from a richer family (SES index = +1.57 versus -0.96); and do twice as much homework (2-3 hours versus less than 1 hour).
[Figure 5 here]
Second, and as previously found for the levels 0 and 1 comparison, the level 1 student with $\Phi^{\text{max}}_{i}$ comes from a small Indigenous rural school (as rate for the nolangtest is equal to 1 and urban = 0), while the student with the highest chance of reaching level 2, comes from a large public school (without language minority students) which is located in an urban (large town) area and with parents involvement fostered by school (index = +0.047) where the latter is not encouraged in the Indigenous school of the level 1 student (index = -0.63) (see Figures 5c and 5d). Level 2 student school’s rates for certification and PD attendance of teachers are high (between 80-95 percent) for the school of student reaching level 2 and they are leading SHAP values, but this is not the case for the other student’s school.
5.4 Countries’ results
To evaluate the possibility of heterogeneity of estimates of Sections 5.2 and 5.3, in this section I by examine SHAP values at the country level. Specifically, I run a similar analysis per country for the level 0 and level 1 sample8 yielding local explanations of SHAP values (students at extremes of the SHAP values distribution). But, first, I present the global contributions—average SHAP (absolute) values for all covariates.9
As a summary, I plot the frequency of the highest SHAP values contributions on predictions across the 10 countries. In Figure 6 I select the top 10 ranked covariates in terms of their SHAP average (absolute) values for each country and then obtain the number of times a given covariate appears in the top 10 list. Figure 6 shows that primary repetition appears in the top ranked list for all 10 countries; while family SES and school climate (in 9 countries) and digital devices (in 8 countries). This is followed by lack of school safety and homework intensity (present in 6 countries), whereas the covariates paid work, mother education and number of books appears in the top ten ranked covariates for 5 countries.
[Figure 6 here]
As regards to local explanations (estimates shown in Tables 6 and 7) —where I run a full model with all the explanatory variables—, the main takeaways concerning leading drivers for students more likely to reach either level 0 or level 1 achievement per country are outlined below.10 For the first set of five countries (Table 6):
  • Argentina. (a) The student with the largest probability of reaching just level 0 (high-index $\mapsto$ level 0) is a repeater, has no devices, comes from a small size school with low parental involvement in the school running and bad school climate. (b) The student with the highest probability of reaching level 1 (low-index $\mapsto$ level 1) comes from better-off family (with positive SES index), has 4 books (2.5 books above mean), attend a private school with just 12% rate of disadvantaged students, and has 11 devices (3 above the mean country value for this sample).
  • Brazil. (a) Profile of student with achievement $\mapsto$ level 0: attend a school with full disadvantage, comes from low SES poor family, has zero books and, surprisingly, an educated father. (b) Profile of student with achievement $\mapsto$ level 1: does more regularly homework (above average), attend a school with below average index of bad school climate, satisfaction with life in the middle of the scale, a low educated mother and most likely to work for paid everyday.
  • Chile. (a) Profile of student with achievement $\mapsto$ level 0: is a repeater who also skipped school (1 whole day last 2 weeks), whose mother only completed lower secondary, attends a school of a size in the bottom quartile with good safety in its surroundings areas. (b) Profile of student with achievement $\mapsto$ level 1: attend a school located in a even safer area and with a even better school climate and with a relatively high contextual shortcomings, has only one sibling, and did not repeat primary.
  • Colombia. (a) Profile of student with achievement $\mapsto$ level 0: comes from a poor family (SES into the bottom half of distribution), has none digital devices and repeated primary, attend a school with quite large index for bad school climate, and whose mother is highly educated. (b) Profile of student with achievement $\mapsto$ level 1: comes from a better-off family and does not engage in paid-work, has 3 books (top 90%), and attend a school without disadvantaged students and with a high degree of autonomy (index large and positive).
  • Dominican Rep. (a) Profile of student with achievement $\mapsto$ level 0: repeated primary, attend a school with poor school climate, comes from low-SES family, is satisfied with life, and in the school there is a support of inclusivity. (b) Profile of student with achievement $\mapsto$ level 1: comes also from a school with poor school climate but with the school located in safer area, the extent of support of inclusivity in the school is larger, does not work for pay and is satisfied with life.
[Tables 6 here]
For the second set of five countries (Table 7), main findings are:
  • Mexico. (a) Profile of student with achievement $\mapsto$ level 0: is engaged fully in paid work, does not have any books, missed school (3 months in a row), and repeated primary. (b) Profile of student with achievement $\mapsto$ level 1: comes from relatively poor family with low SES index, but is not engaged in paid work and does regularly homework, attend a school located in an unsafe area, and in a school with parents have low participation.
  • Panama. (a) Profile of student with achievement $\mapsto$ level 0: attend a highly disadvantaged school with a lack of PCs connected to the internet, does hardly any homework, comes from really poor household whose mother did not complete ISCED level 1 (primary). (b) Profile of student with achievement $\mapsto$ level 1: comes from a family whose SES is in the top quartile and has 9 digital devices (large than the median by 2), though the school he//she attends has very low autonomy and no qualified teachers.
  • Peru. (a) Profile of student with achievement $\mapsto$ level 0: comes from an Indigenous background who repeated primary and does just less than 1 hour homework per week, and attends a school with severe low school climate and very weak ICT infrastructure. (b) Profile of student with achievement $\mapsto$ level 1: does not work for payment, has 8 digital devices at home (above median), the family wealth falls into the top half of the SES distribution, an attend school with a STR of 32 with high autonomy.
  • Paraguay. (a) Profile of student with achievement $\mapsto$ level 0: comes from a household with 3 devices and high SES, only engage 1 day in unpaid work, but attend a school with low autonomy and just 16% of teacher attending PD. (b) Profile of student with achievement $\mapsto$ level 1: is richer than a level 0 student (higher value of SES index) and has more devices at home, attend a private school with larger autonomy of a total size of $\approx$ 900 students.
  • Uruguay. (a) Profile of student with achievement $\mapsto$ level 0: primary repeater, low life satisfaction, no devices at home, whereas (b) Profile of student with achievement $\mapsto$ level 1: did not repeat primary, family SES is larger (in the 3rdrd quartile), and has 3 books at home (median for this data is just 1), attends a school with more than 70% of either certified teachers or teachers attending PD.
[Tables 7 here]
6 Discussion
Latin American students’ limited achievement at secondary level, underpinned by systemic embedded inequalities, are very concerning. This inequality-driven learning crisis has been made worse by the COVID-19 pandemic (Betthäuser et al., 2023; UNESCO, 2024), with the region experiencing the larger period of school closures (of around 269 days, Bracco et al., 2025; Gómez and Andrés Uzín P, 2022). The region learning crisis is severe, with only between 5 and 7 out of 10 students reaching adequate competency in reading and math (Arias Ortiz et al., 2024b). Is within this bleak scenario that the paper’s contribution should be understood, namely, the profiling of low achievers, so that adequate interventions can ameliorate the extent of the learning crisis in the region.
In particular, I follow a mechanism to identify poor performing students through explainable ML methods (Chen et al., 2022; Lundberg and Lee, 2017; Štrumbelj and Kononenko, 2014), i.e., Shapley values. The paper contributes to the emerging body of studies using interpretable ML results in education (Ali et al., 2025; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022; Zhu et al., 2025), though chiefly, expanding this literature to global south education systems at a regional scale.
The paper, relying equally in global and local explanations at a regional level and per country, yields several insights into the LAC region learning crisis by offering a detailed account of which factors may sturdily place a students in the lowest bracket of performance (level 0, no competency achieved) or, alternative, which malleable factors may allow a bottom performer to transition to the next achievement bracket, level 1. The same analysis is carried out for the level 1 to level 2 students (where in level 2 basic subjects competencies are achieved). This granular analysis at the bottom of the learning distribution is fundamental for successful targeted interventions pushing up students to the next level of achievement. In essence, the paper’s approach and its policy implications are close to the existing literature of the region for the detection of secondary students at risk of dropout (e.g., Haimovich and Adelman, 2021; Rodríguez et al., 2023; Vinas-Forcade et al., 2021) via the setup of automatic alert systems (Early Warming Systems).
The utility of Local SHAP values cannot be undervalued. They are a powerful policy tool as they offer an identification strategy of students locked into lower performing groups (or at risk of falling into that group) because they, not only rank leading covariates revealing positive or negative impacts of each variable on the model output for each observation (student), but also because they give the specific values each covariates take in that ranking, thereby given precise profiles of students. Results from the SHAP method can further guide follow-up qualitative studies attempting to gather more in-depth information on the reasons behind poor performance, a sequential combination of ML approaches and interviews used in late research (Gonzalez, 2021; Jayachandran et al., 2023). The sequential application of estimated SHAP values in learning surveys with qualitative interviews (mixed methods), is an interesting avenue for future research which can offer more insights into the learning crisis problem. Below, I flesh out policy implications from leading regional estimates which can be framed as a gradual two-way process of interventions: (i) first, to identify those at the bottom of achievement; and (ii) to introduce leverage mechanisms to modify covariates of level 0 (or level 1) students towards profiles of level 1 (level 2) students.
In the local regional analysis, I find that the profile for a student who is locked-into the lack of competency group (level 0) is Indigenous, a repeater, is engaged in paid work (half of the week) and comes from a very poor household without digital devices. Thus, an entry point of a policy is to target poor Indigenous students and//or poor repeaters with financial support so that to avoid (or minimise) the impact on learning of paid work, likely leading to repetition. Culture responsive programs on bilingual education, raising Indigenous agency (López, 2021) and the attraction of good teachers to schools with mostly Indigenous students (Cortina, 2017; Delprato et al., 2022), tackling also repetition earlier on in primary (see: Learning Spaces program in Dominican Rep; Chávez et al., 2021), and conditional cash transfers (e.g., see examples of Brazil and Peru Draeger, 2021; Gaentzsch, 2020) alleviating disadvantaged students financial constraints, would all be beneficial interventions for level 0 students.
Conversely, I find that students performing at level 1 do not work, are not repeaters, have a higher number of books and digital devices at home. Hence, shifting leading covariates from the profile of level 0 performing student to a level 1 performing student, would also require additional public investment on educational inputs. This shift, in turn, is related to programs such as the one laptop books in the region (e.g in rural schools in Peru: Cristia et al., 2017, or in Uruguay: Díaz et al., 2022; Melo et al., 2017). Note as well that, improving school climate and boosting students’ well-being support from teachers, are features of level 1 student profile which can benefit the lowest level 0 performers. The overall importance of school climate for learning is in line with the recent reviews of Varela et al. (2024) and Chávez et al. (2021), and the need of targeting boys of public schools in the region as an initial step (Delprato et al., 2017).
Besides, a comparison of school features between these two level 0 and level 1 students’ profiles —and as found by earlier studies (Agasisti et al., 2023; Delprato, 2025; Mereles and Canese, 2022)— highlights the need of higher investments in schools’ ICT stock and connectivity, and promoting the collaboration on the use of digital devices by teachers (Arias Ortiz et al., 2024a). As shown by earlier literature, weak poor ICT access in the LAC region is indeed a barrier for those students attempting to move up in the learning distribution towards performance levels where subjects competencies are in fact reached. For instance, CEPAL (2020)’s report indicates that only 70% of people in the LAC region poses internet connection, and just 46% of children (aged 5-12) live in households without an internet connection. High gaps on ICT access by socioeconomic status are also observed, with students in the top wealth quartile being at least four times as likely (in comparison to the students from the bottom quartile) to have a laptop at home (Berlanga et al., 2020). This comparison also reveals that more investment is required on teachers preparation and teachers’ numbers in pedagogic roles (Ainscow et al., 2025; Miranda et al., 2025), ensuring quality and promoting continuous professionalisation.
Similarly, policies seeking out to move up students from level 1 to level 2 would need to focus on —as well as enhancing level 1’s students stock of educational inputs and alleviating poverty and the chances of being involved in paid work—, offering better conditions for level 1 students to be able to carry out more homework (e.g., after school classes), and boosting family educational support. At the school level, policies should reshape schools’ functioning on different dimensions: boost involvement on parents on schools’ governance, and to enhance teaching quality too (via attracting teachers having more certification and attendance of PD training).
The analysis has some limitations. Methodologically, a more exhaustive approach for benchmarking students profiles would be helpful, producing a more robust contrasting of students’ profiles, based on incremental steps (e.g., deciles) of the SHAP values index $\big[\Phi^{\text{min}}_{i}$, $\Phi^{\text{q10}}_{i}$$\Phi^{\text{q90}}_{i}$, $\Phi^{\text{max}}_{i}\big]$. Rather than focusing on two students for each sample, a continuum of profiles along students’ distribution of SHAP values would allow a better comparison on shift of covariates underpinned by future scaffolding interventions. Also, the use of SHAP values has found some methodological critics (Kumar et al., 2020), though this critic is an overall problem for model explanation methods (Chen et al., 2023). A future direction of research is to address intersectionality (combinations of student//family covariates with school covariates) which is a driving concept behind how inequality appears and persist in resource-constrained education systems (Codiroli Mcmaster and Cook, 2019). I succinctly address this in Appendix E (Figure E1) with SHAP values interactions, though further work in this direction is needed so that interventions are designed using joint markers of personal and contextual disadvantages from students and schools.
7 Conclusions
In this paper, I analyse the determinants of low performers students in the Latin American region at secondary level using PISA 2022 data, including 10 LAC countries. Combining the performance on three subjects assessed (math, reading and science), I focus on the bottom learning distribution into the three brackets of performance (OECD, 2023a): level 0 (no achievement), level 1 (below basic competencies) and level 2 (achieving basic competencies), where around 83% of the region students’ performance falls. I rely on two working samples: the sub-groups of bottom performers (level 0 versus level 1, $N$ = 16,236) and low performers (level 1 versus level 2, $N$ = 9,484) to compare the dealing determinants of only reaching either level 0 or level 1, correspondingly. Methodologically, using recent advancements on interpretable ML modelling (Chen et al., 2023; Lundberg and Lee, 2017), I use the SHAP method to pin down the main determinants of learning poverty in the LAC region. I obtain global and local versions of estimates, both at the regional level and country level.
The main findings from the global explanatory analysis can be summarised as follows. First, variables with the highest impacts (regardless of the direction of associations) are, to some extent, similar for determinants for bottom and low performers. Primary repetition, sense of belonging to school and well-being support from teachers are key drivers of falling into the lowest achievement group, whereas student’s paid work, the intensity of homework and mother’s educational level are relevant for the level 1 versus level 2 comparison (located higher in the learning distribution), with household wealth and educational support and digital devices being higher ranked than in the level 0-level 1 comparison. Second, the rate of school contextual poverty and STR (pedagogic support) are also more relevant for the reaching level 2, whilst the prevalence of students with a second language being more prominent as a driver for the level 0 performance group.
Third, local estimations for students placed at the extremes of SHAP index distribution, indicate that a student with the highest probability of being a level 0 achiever attends a school with bad school climate, speak other minority language and had repeated, has no digital devices at home, comes from a poor family and works for payment half of the week. The school profile of this student shows wide disadvantages such as weak ICT infrastructure and teaching quality (only a third of teachers being certified), whereas a level 1 achiever attends a private with more autonomy and a low STR for pedagogic staff, more qualified teachers, and with larger connectivity (over 90% of PCs connected to internet). Similarly findings are obtained for the level 1 and level 2 comparison where —as well as a level 2 student being better off financially overall, and having more educational inputs and support—, he//she is not in the labour market having more feasible time for doing homework (twice as much as a level 1 student), and attends a school with considerable higher school’s rates for certification and PD attendance of teachers (80%-95%).
Regarding countries’ estimates, I find quite homogenous patterns as far as global average contribution of top ranked factors is concerned, with repetition at primary, household wealth, and educational ICT inputs (digital devices) being top ten ranked covariates in at least 8 out of the 10 total countries. Other important covariates across countries are school safety, student work, books at home and mother’s education. There is, however, some between-country heterogeneity when it comes to local estimates and defining characteristics of a lowest achieving student. Level 0 (no achiever) shares, across countries, features such as being a repeater with very few devices, coming from a low SES family and attending a school with poor school climate, but with some differences and specific country drivers. In Chile, the lowest performer also skipped school and the education of his//her mother is below secondary while in Panama mothers’ education for this level 0 student profile is even worse (uncompleted primary); in Mexico, a new feature is that a level 0 student does paid-work for nearly the whole week (and missed school for 3 months); in Peru and Paraguay school features turn out to be prominent (very low ICT infrastructure, low autonomy and just 16% of PD teachers attendance in the case of Paraguay, and the lowest performing student for Peru is Indigenous). A Level 1 student profile is, in the majority countries, from a family with an above average level of wealth, has a fair number of digital devices, is not engaged in paid work, and mostly their schools they attend have a large degree of autonomy on their decisions.
Notes
Appendix 0.A Evaluation metrics
Here I describe the different metrics to assess the five ML models used with a distinct hyperparameters configurations. The Accuracy (ACC) metric equation is as follows:
[…]
where TP (TN) is the true (negative) positive rate, and FP (FN) is the false positive (negative) rate. The second metric, recall (RC), measures the capacity of the model to efficiently identify all positive cases, and it is defined as:
$RC=\frac{\text{TP}}{\text{TP + FN}}$ (7)
The precision metric (PR) quantifies, among all instances of classified positive responses, those accurately identified as positive responses:
$PR=\frac{\text{TP}}{\text{TP + FP}}$ (8)
The complement of the PR metric is the specificity metric (SP) which measures the percentage of correctly classified negative samples (true negatives) with respect to all samples classified as negative, and it is given by:
$SP=\frac{\text{TN}}{\text{TN + FP}}$ (9)
The F1 score metric combines the precision and recall scores of a model by using the harmonic mean of PR and RC, yielding a balance between the model accuracy and its completeness. F1 is obtained as:
$ACC=\frac{2\times(\text{PR}\times\text{RC})}{\text{PR + RC}}$ (10)
Finally, the Area Under the Curve (AUC) metric is defined as the integral of the ROC curve, allowing a assessment of the classification performance for varying discrimination thresholds, with a AUC equals to one as perfect classification and a AUC of 0.5 equals to a random classification. The AUC is calculated as follows:
$AUC=\int_{0}^{1}ROC(t)dt$ (11)
Appendix 0.B Shapley values and ML implementation
[…]
Shap values—estimation. When estimating a ML model, the aim is to explain a predictive model $f(x)$ trained on $\mathcal{X}=\{x^{[i]},y^{[i]}\}_{i=1}^{N_{\text{train}}}$, where $x^{[i]}$ is a feature vector of dimension $M$, $y^{[i]}$ is the vector of responses and $N_{\text{train}}$ is the number of observations used in the trained model. The prediction for a given feature vector value ($\widehat{y}=f(x=x^{*}$)) is obtained using Eq. (12) where the predictive model $f$ is the game, and the feature vector is the $M$ players. The $j$th feature in the prediction $f(x^{*})$ is given by the Shapley value $\phi_{j}$, where $f(x^{*})=\phi_{0}+\sum_{j=1}^{M}\phi^{*}_{j}$ and so the sum of Shapley values is the difference between the given prediction and the global average prediction (i.e., $\phi_{0}=\mathbb{E}[f(x)]$). To calculate Eq. (12), one needs to define a function which mirrors $f(x^{*})$ when only features on the coalition $\mathcal{S}$ are known: $\upsilon(\mathcal{S})=\upsilon(\mathcal{S},x^{*})$. This can be approximated by the contribution function (the expected response $f(x)$ conditioning on the features in $\mathcal{S}$ taking on the values $x_{\mathcal{S}}^{*}$):
$\upsilon(\mathcal{S})=\mathbb{E}\big[f(x_{\mathcal{\bar{S}}},x_{\mathcal{S}}|x_{\mathcal{S}}=x^{*}_{\mathcal{S}})\big]$ (13)
where $x_{\mathcal{\bar{S}}}$ denotes features outside the subset $\mathcal{S}$. The conditional expectation in Eq. (13) describes the whole probability distribution, and is a common estimator for predictions’ applications. Here, for the estimations of leading drivers of learning poverty, I follow Lundberg and Lee (2017) who introduced the SHAP (SHapley Additive exPlanation) method which unifies the class of additive feature attribution methods. There are various (interventional conditional expectation) algorithms to define the value function for a set of attributes (e.g., DeepSHAP, LinearSHAP, KernelSHAP, TreeSHAP).
In this paper, and due to computational issues and the fact that I use a stacked ML model (which is a combination of classification models), I rely on the LinearSHAP (Chen et al., 2020), the default Explainer of the SHAP Python package, which estimates the conditional expectation of $x$ given different subsets (for a model $f(x)=\beta X+b)$:
$\phi_{i}(f,x)$ $=$ $\frac{1}{M!}\sum_{R}\mathbb{E}\big[f(x)|x_{\mathcal{S}^{R}\cup i}\big]-\mathbb{E}\big[f(x)|x_{\mathcal{S}^{R}}\big]$ (14)
$=$ $\beta\frac{1}{M!}\sum_{R}\mathbb{E}\big[x|x_{\mathcal{S}^{R}\cup i}\big]-\mathbb{E}\big[x|x_{\mathcal{S}^{R}}\big]$ (15)
A key feature for the estimations which are used in the paper are SHAP feature importance plots, offering a summary of the contribution of each feature over a set of samples. The importance of each feature is calculated by taking the average of the absolute Shapley values per feature across all samples, using the following formula:
[…]
Table C1: Students and family variables description
Variable Name Question Original variable Variable type Variable label
(1) (2) (3) (4) (5)
Gender – female sex_f Q3 st004d01t Binary 1=female, 0=male
Satisfaction with life satislife Q42 st016q01na Discrete 1=lowest,…,10=highest
Homework – all subjects (time spent) homework Q44 st296q04ja Discrete 1 (< 30min), 2 (30min-1hour), 3 (1- 2 hours),
4 (2-3 hours), 5 (3-4 hours), 6 (> 4 hours)
Digital devices at home-number devices Q8 st254q02ja to st254q06ja Discrete 1 to 20
Books – number books Q9 st255q01ja Discrete 0 (none), 1 (1 to 10), 2 (11 to 25), 3 (26 to100),
4 (101 to 200), 5 (201 to 500), 6 (> 500)
Immigrant immigrant_yes Q20 st021q01ta Binary 1=immigrant, 0=no immigrant
Speak other language at home otherlang_yes Q21 st022q01ta Binary 1=speak other language at home, 0=otherwise
Late entry – primary lateprim_yes Q24 st126q01ta Binary 1=late entrant in primary, 0=otherwise
Repetition – primary repeprim_yes Q25 st127q01ta Binary 1=repeated primary, 0=otherwise
Repetition – lower secondary repelowsec_yes Q25 st127q02ta Binary 1=repeated lower secondary, 0=otherwise
Missed school (3 months in a row) missclass_yes Q26 st260q02ja Binary 1=missed school 3 months in a row (lower secondary), 0=otherwise
Skipped school (1 day last 2 weeks) skippsch_yes Q28 st062q01ta Binary 1=skipped school (1 whole day last 2 weeks), 0=otherwise
Teacher support for well being (index) tchsuppwb Q29 st267q01ja to st267q08ja Continuous Intervals: low (-4.634, 0.053) and high (0.054, -14.024)
Sense of belonging to school (index) belongsch Q30 St034q01ta to st034q06ta Continuous Intervals: low (-5.110, -0.039) and high (-0.038, 4.326)
School climate – bad (index) schclimabad Q31 st038q03na to st038q11ja Continuous Intervals: low (-0.753, -0.296) and high (-0.297, 9.336)
Safety in school and surroundings (index) schsafelack Q32 st265q01ja to st265q04ja Continuous Intervals: low (-1.257, 0.357) and high (0.358, 4.077)
Work – unpaid, days workunpaid Q34 st294q03ja Discrete 0 (0 day), 1 (1 day), 2 (2 days), 3 (3 days), 4 (4 days), 5 ( $ge$ 5 days)
Work – paid, days workpaid Q35 st294q04ja Discrete As above
Mother education motheredu Q11 st005q01ja Discrete 1 (<< ISCED1), 2 (ISCED level 1), 3 (ISCED level 2), 4 (ISCED level 3.3),
5 (ISCED level 3.4), 6 (ISCED level 4),…,8 (ISCED level 8)
Father education fatheredu Q13 st007q01ja Discrete As above
Siblings number siblings Q10 st230q01ja Discrete 1 (one), 2 (two), 3 (3 or more)
SES (index) ses derived na Continuous Intervals: low (-6.654, -0.852), high (-0.851, 2.039)
Family support for education (index) famsuppedu Q55 st300q04ja to st300q10ja Continuous Intervals: low (-2.728, 0.174), high (0.174, 1.185)
  • (1) For further details on student and family variables’ definition and measurement, check PISA 2022 Student Questionnaire: link.
Table C2: School variables—description
Variable Name Question Original variable Variable type Variable label
(1) (2) (3) (4) (5)
Urban urban Q1 sc001q01ta Binary 1=urban, 0=rural
Private private Q2 sc013q01ta Binary 1=private, 0=public
Town size townsize Q1 sc001q01ta Discrete 1 (village: < 3000), 2 (small town: 3000-15000),
3 (town: 15000-100000), 4 (city, large city: \$ge\$ 1million)
Competition competition Q5 sc011q01ta Binary 1=competition, 0=no competition
Funding government fundgovt Q4 sc016q01ta Continuous 0-1 (rate)
Funding fees fundfees Q4 sc016q02ta Continuous 0-1 (rate)
Students – second language nolangtest Q7 sc211q01ja Continuous 0-1 (rate)
Students - learning needs learnneeds Q8 sc211q02ja Continuous 0-1 (rate)
Students – disadvantaged disadvantaged Q9 sc211q03ja Continuous 0-1 (rate)
School size (enrolment) enrolment Q6 sc002q01ta to sc002q02ta Continuous 0 to 7000
STR srt Q6,Q8 sc018q01ta01, sc018q01ta02 Continuous 0 to720
Teachers attended PD programme (last 3 months) tchpdatt Q21 sc025q01na Continuous 0-1 (rate)
STR non-teaching staff, pedagogic support strpedag Q6, Q10 sc168q01ja Continuous 0-1 (rate)
Teachers – certified tchcertified Q8 Sc018q02ta01, sc018q02ta02 Continuous 0-1 (rate)
Teachers – degree or higher tchdegree Q8 sc018q08ja01 to sc018q09ja02 Continuous 0-1 (rate)
ICT – infrastructure per student (index) ictinfra Q15 sc004q01ta,sc004q02ta, sc004q03ta, Continuous Intervals: low (-4.683, -0.060) and high (0.061, 37.96)
sc004q07na, sc004q08ja
PC connected web-teacher ratio pcwebtr Q15 Sc004q01ta, sc004q03ta Continuous 0-1 (rate)
Admission criteria admissionnone_yes Q11 sc012q01ta to sc012q12ja Binary 1=admission (any criteria used), 0=otherwise
Transfer policies transfernone_yes Q12 sc185q01wa sc185q03wa sc185q04wa Binary 1=transfer (due to performance and behaviour), 0=otherwise
Autonomy (index) autonomy Q13 sc202q01ja to sc202q12ja Continuous Intervals: low (-2.268, 0.068) and high (0.069, 1.794)
Social and emotional skills policy polsocemskill_yes Q48 sc189q01ja Binary 1=policy exists, 0=otherwise
Teacher monitoring – methods tchmonitor Q19 sc032q01ta to sc032q04ta Discrete 0, 1, 2, 3 and 4
Testing students learning, all methods teststud_yes Q37 sc034q01na sc034q02na sc034q03ta Binary 1=test, 0=otherwise
Staff students support for inclusivity (index) suppinclusity Q28 sc173q01ja to sc173q06ja Continuous Intervals: low (-2.290, 0.084) and high (0.085, 1.399)
Parents involvement fostered by school (index) parentinvol Q30 sc192q01ja to sc192q06ja Continuous Intervals: low (-2.666, -0.043) and high (-0.042, 2.490)
School policy – ability grouping abilitygroup_yes Q39 sc042q01ta, sc042q02ta Binary 1=ability grouping, 0=otherwise
  • (1) For further details on school variables’ definition and measurement, check PISA 2022 School Questionnaire: link.
Table D1: Outcome (performance level 0 versus level 1). Comparison of ML models – parameters search
Model Parameters Evaluation metrics
ACC RC F1 PR SP AUC
(1) (2) (3) (4) (5) (6)
Logistic penalty = none 0.7305 0.8609 0.7591 0.4846 0.8068 0.7825
regression penalty = l2 0.7299 0.8548 0.7613 0.4943 0.8053 0.7823
Lasso C = 10 0.7533 0.8641 0.7816 0.5443 0.8208 0.8143
with CV penalty C = 50 0.7549 0.8702 0.7801 0.5373 0.8227 0.8139
C = 100 0.7527 0.8576 0.7843 0.5549 0.8193 0.8147
C = 200 0.7524 0.8571 0.7842 0.5549 0.8190 0.8147
Gradient boosting Trees = 100, learning rate = 0.01 0.6989 0.9725 0.6918 0.1826 0.8085 0.7871
Trees = 100, learning rate = 0.10 0.7752 0.8730 0.8010 0.5909 0.8354 0.8560
Trees = 500, learning rate = 0.01 0.7588 0.8813 0.7788 0.5277 0.8269 0.8378
Trees = 500, learning rate = 0.10 0.7737 0.8632 0.8048 0.6049 0.8330 0.9037
Trees = 1000, learning rate = 0.01 0.7752 0.8748 0.8000 0.5874 0.8357 0.8559
Trees = 1000, learning rate = 0.10 0.7755 0.8553 0.8115 0.6251 0.8328 0.9292
Random forest Trees = 100, max depth = 5 0.7503 0.8474 0.7869 0.5672 0.8160 0.8074
Trees = 100, max depth = 10 0.7616 0.8571 0.7943 0.5812 0.8245 0.9305
Trees = 500, max depth = 5 0.7533 0.8483 0.7899 0.5742 0.8180 0.8083
Trees = 500, max depth = 10 0.7673 0.8609 0.7988 0.5909 0.8287 0.9319
Trees = 1000, max depth = 5 0.7530 0.8492 0.7890 0.5716 0.8180 0.8086
Trees = 1000, max depth = 10 0.7692 0.8627 0.7998 0.5926 0.8301 0.9330
Neural net 1 hidden layer (200 nodes), activation = relu 0.6676 0.6598 0.7966 0.6822 0.7218 0.7565
1 hidden layer (200 nodes), activation = logistic 0.7488 0.8395 0.7895 0.5777 0.8137 0.8704
1 hidden layer (200 nodes), activation = tanh 0.7296 0.7604 0.8137 0.6716 0.7861 0.8573
3 hidden layers (200, 100 and 50 nodes), activation = relu 0.7302 0.8432 0.7671 0.5171 0.8034 0.8676
3 hidden layers (200, 100 and 50 nodes), activation = logistic 0.7479 0.8423 0.7870 0.5698 0.8137 0.8773
3 hidden layers (200, 100 and 50 nodes), activation = tanh 0.7403 0.8399 0.7797 0.5522 0.8087 0.8599
  • (1) Within each learner//model, highest metrics are shown in bold. (2) Also, the chosen most efficient configuration of each model is highlighted in bold –that is, models with the largest number of top preforming metrics– and in case of a draw, I rely on the AUC metric (column 6) to reach a final selection.
Table D2: Outcome (performance level 0 versus level 1). Comparison of ML models – parameters search
Model Parameters Evaluation metrics
ACC RC F1 PR SP AUC
(1) (2) (3) (4) (5) (6)
Logistic penalty = none 0.7305 0.8609 0.7591 0.4846 0.8068 0.7825
regression penalty = l2 0.7299 0.8548 0.7613 0.4943 0.8053 0.7823
Lasso C = 10 0.7533 0.8641 0.7816 0.5443 0.8208 0.8143
with CV penalty C = 50 0.7549 0.8702 0.7801 0.5373 0.8227 0.8139
C = 100 0.7527 0.8576 0.7843 0.5549 0.8193 0.8147
C = 200 0.7524 0.8571 0.7842 0.5549 0.8190 0.8147
Gradient boosting Trees = 100, learning rate = 0.01 0.6989 0.9725 0.6918 0.1826 0.8085 0.7871
Trees = 100, learning rate = 0.10 0.7752 0.8730 0.8010 0.5909 0.8354 0.8560
Trees = 500, learning rate = 0.01 0.7588 0.8813 0.7788 0.5277 0.8269 0.8378
Trees = 500, learning rate = 0.10 0.7737 0.8632 0.8048 0.6049 0.8330 0.9037
Trees = 1000, learning rate = 0.01 0.7752 0.8748 0.8000 0.5874 0.8357 0.8559
Trees = 1000, learning rate = 0.10 0.7755 0.8553 0.8115 0.6251 0.8328 0.9292
Random forest Trees = 100, max depth = 5 0.7503 0.8474 0.7869 0.5672 0.8160 0.8074
Trees = 100, max depth = 10 0.7616 0.8571 0.7943 0.5812 0.8245 0.9305
Trees = 500, max depth = 5 0.7533 0.8483 0.7899 0.5742 0.8180 0.8083
Trees = 500, max depth = 10 0.7673 0.8609 0.7988 0.5909 0.8287 0.9319
Trees = 1000, max depth = 5 0.7530 0.8492 0.7890 0.5716 0.8180 0.8086
Trees = 1000, max depth = 10 0.7692 0.8627 0.7998 0.5926 0.8301 0.9330
Neural net 1 hidden layer (200 nodes), activation = relu 0.6676 0.6598 0.7966 0.6822 0.7218 0.7565
1 hidden layer (200 nodes), activation = logistic 0.7488 0.8395 0.7895 0.5777 0.8137 0.8704
1 hidden layer (200 nodes), activation = tanh 0.7296 0.7604 0.8137 0.6716 0.7861 0.8573
3 hidden layers (200, 100 and 50 nodes), activation = relu 0.7302 0.8432 0.7671 0.5171 0.8034 0.8676
3 hidden layers (200, 100 and 50 nodes), activation = logistic 0.7479 0.8423 0.7870 0.5698 0.8137 0.8773
3 hidden layers (200, 100 and 50 nodes), activation = tanh 0.7403 0.8399 0.7797 0.5522 0.8087 0.8599
  • (1) Within each learner//model, highest metrics are shown in bold. (2) Also, the chosen most efficient configuration of each model is highlighted in bold –that is, models with the largest number of top preforming metrics– and in case of a draw, I rely on the AUC metric (column 6) to reach a final selection.
Figure E1: Intersectionality—SHAP values interactions
Notes: (1) The heatmap plots of correlations across SHAP values include the top 20 covariates across students//family and school features. (2) Top plot (level 0 versus level 1) and bottom plot (level 1 versus level 2).
Figure E2: Model: level 1 versus level 2 (low performing group). Beeswarm dot plot and heatmap of individual SHAP values for each student//famly and school covariate across observations
Notes: (1) Beeswarm plot. Red represents that covariates is = 1 and, blue (= 0) otherwise. (2) Heatmaps. Colour intensity reflects the degree of impact on student performance, ranging from low (blue) to high (red).
(a) Argentina
(b) Brazil
(c) Chile
(d) Colombia
(e) Dominican Rep.
(f) Mexico
(g) Panama
(h) Peru
(i) Paraguay
(j) Uruguay
Figure F1: Countries’ results for model: level 0 versus level 1 (bottom performing group). Student//famliy and school covariates average SHAP values (absolute values)